# Dyna-Think: Synergizing Reasoning, Acting, and World Model Simulation in AI Agents

### v1

@@ 4.2 Experimental Setup
Table 1: Number of tasks per domain in the training and test set.
Domain Train Test
OS 57 31
VSCode 46 38
Chrome 88 54
GIMP - 32
TBird - 19
Since most computer-use benchmarks were not designed for training, the number of tasks available in each domain is often limited. We thus construct a training and test set by 1) manually augmenting existing tasks in each domain to increase its size; 2) construct training/test splits for each domain, and 3) held out two domains (GIMP and Thunderbird) from training to separately measure In-domain (ID) and Out-of-domain (OOD) performance. To augment a task, we follow the principle that 1) action sequence that correctly completes task $A$ does not complete the augmented task $A^{\prime}$; and 2) the augmented task $A^{\prime}$ can still be evaluated using OSWorld’s evaluation scripts after some adjustments. Please refer to Section E.2 for more details and example augmented tasks. In Table 1 we report the training and testing dataset statistics. To promote reproducibility, we will release these datasets along with our codebase.
@@
We evaluate all runs using the accessibility-tree mode (i.e., text-only), and report task success for ID and OOD tasks. We use the same inference prompts and hyperparameters (e.g., temperature of 1.0) provided by OSWorld, but extend the maximum number of steps per task to 30 in order to measure scaling abilities. To provide a more robust evaluation for these long-horizon tasks, we report the average success rate (Avg) and the Best-of-N success rate (BoN) over 3 runs. Each evaluation is ran with 3 threads, lasting on average 18-24 hours to complete.
@@ 4.3 Main Results
We compare our Dyna-Think framework against 1) training-free methods based on expert LLMs; and 2) related training methods. For training-free methods, we consider directly prompting LLMs such as o3-mini and DeepSeek-R1. For training-based methods, we consider 1) Reinforcement Finetuning (RFT) which only peforms policy learning, by finetuning on correct rollouts after rejection sampling (Best-of-3); and 2) the vanilla Dyna algorithm (we follow Fang et al. (2025)) which performs world model learning with a separate language model $\mathcal{W}(\mu)$, and then trains the policy using correct rollouts [^1] When rolling out with the learned $\mathcal{W}(\mu)$, correctness of a trajectory is evaluated using an LLM (Fang et al., 2025). We use GPT-4o in this work. collected with both the real $\mathcal{W}$ and the learned $\mathcal{W}(\mu)$. Our DDT method only uses rollouts with $\mathcal{W}$ and performs policy and world model learning on a single model. For a fair comparison, we use the same set of rollout trajectories (116 in total) to construct world model (116) and policy training data (35 after rejection sampling, a similar data size as Yu et al. (2024b)) for all training methods, whenever possible.
@@
- (GPT-4o-2024-11-20) - - 1.0x 16.5$\pm$2.5 19.5 9.2 31.6 36.6 19.6
- (Qwen2.5-32B-Instruct) - - 0.2x-0.8x 14.2$\pm$2.2 18.9 2.5 27.0 35.0 7.8
- (R1-distill-70B) - - 0.6x-2.2x 7.6$\pm$2.0 9.8 2.5 15.5 20.3 3.9
- (o3-mini-2025-01-31) - - 3.4x-7.0x 20.5$\pm$1.4 23.3 13.7 31.0 36.6 17.6
- (R1) - - 1.7x-4.9x 31.2$\pm$1.4 35.8 20.2 44.8 48.8 35.3
DIT(R1) - - 1.1x-2.5x 22.6$\pm$1.0 26.8 12.4 35.6 40.7 23.5
   +RFT 35 - 1.0x-2.7x 23.2$\pm$1.0 26.0 16.3 38.5 43.1 27.5
   +vanilla Dyna 94 116 1.1x-2.5x 24.1$\pm$2.1 26.8 17.6 35.6 29.3 31.4
   +DDT($\hat{\mathcal{T}}$) 35 116 1.1x-2.7x 25.9$\pm$1.4 28.7 19.0 43.1 47.2 33.3
   +DDT($\hat{\mathcal{T}}_{\Delta}$) 35 116 1.2x-3.5x 23.2$\pm$0.7 27.6 12.4 38.5 44.7 23.5
   +DDT($\hat{\mathcal{T}}_{\textrm{critic}}$) 35 116 1.2x-2.7x 26.6$\pm$1.5 30.3 17.6 44.3 49.6 31.4
@@
- (GPT-4o-2024-11-20) 1.0x 16.5$\pm$2.5 19.5 9.2 31.6 36.6 19.6
- (Qwen2.5-32B-Instruct) 0.2x-0.8x 14.2$\pm$2.2 18.9 2.5 27.0 35.0 7.8
- (R1) 1.7x-4.9x 31.2$\pm$1.4 35.8 20.2 44.8 48.8 35.3
Direct Distill(4o) 0.5x-1.7x 15.3$\pm$2.4 18.5 7.8 28.7 34.1 15.7
Direct Distill(R1 no-think) 0.1x-0.5x 17.1$\pm$2.1 19.3 11.8 28.2 30.9 21.6
Direct Distill(R1) 1.6x-6.0x 20.9$\pm$1.0 24.6 11.8 36.2 42.3 21.6
DIT(R1) 1.1x-2.5x 22.6$\pm$1.0 26.8 12.4 35.6 40.7 23.5
Results
We report our results in Table 2. We first compare training-based methods. In Table 2, we find that DDT training, especially when trained with next-state prediction (DDT($\hat{\mathcal{T}}$)) and critique prediction (DDT($\hat{\mathcal{T}}_{\textrm{critic}}$)) improves upon both RFT and vanilla Dyna in both average success rate and BoN success rate. This indicates that 1) world model training benefits policy training; and 2) directly training $\pi_{\mathcal{W}}(\theta)$ as a world model is more effective than planning with a separately trained $\mathcal{W}(\mu)$. Next, we find that training with critique data showed strong performance compared to training on state-difference (DDT($\hat{\mathcal{T}}_{\Delta}$)) and on next-state prediction (DDT($\hat{\mathcal{T}}$)). We believe this is because these critique data provide a more interpretable signal for $\pi_{\mathcal{W}}(\theta)$ to improve its world modeling and planning ability during inference (see Section 5.3 for more quantitative study). Moreover, we find that training with the next-state prediction objective (DDT($\hat{\mathcal{T}}$)) shows strong performance for OOD tasks. We believe this is because next-state prediction more effectively enhanced the model’s understanding of the environment, which is useful for OOD tasks where many states are entirely novel. Finally, compared to best training-free methods such as prompting R1(685B), we find Dyna-Think can achieve similar BoN performance, while generating 2x less tokens on average and being only 32B in size. These results indicate the effectiveness of Dyna-Think at improving agent performance.
@@ 4.4 Thinking Behavior Analysis
We present the results in Table 3. First, we find that Direct Distill(R1 no-think) deteriorates significantly compared to Direct Distill(R1), even though they are trained on the same set of trajectories. This indicates that including long-CoT thinking during policy training is benefitial. Next, we find that DIT(R1) achieved similar performance compared to Direct Distill(R1), despite generating 2x less tokens on average. This suggests that the ability to perform reasoning with world-model simulations is the central part in R1-style thinking, underscoring the importance of world modeling in AI agents.
@@ 5.1 Scaling World Model Training in Dyna-Think
Figure 4: Scaling world model learning with synthetic tasks generated by GPT-4o. We use DDT($\hat{\mathcal{T}}_{\textrm{critic}}$) and train from our best model in Table 3. After world model training, we perform one-round of policy training using the same set of policy learning data.
Since world model learning only requires interaction data (without evaluators for task success) with the real $\mathcal{W}$, we investigate whether scaling world training in Section 3.3 can further improve policy performance. To test this, we 1) prompt GPT-4o to generate synthetic task instructions for each domain; 2) use $\pi_{\mathcal{W}}(\theta)$ collect rollout trajectories; and 3) construct a world model learning dataset following Section 3.3. Since no evaluator is required for world modeling data construction, we used all rollout trajectories that terminated within a maximum of 30 steps. This results in a total of 703 additional trajectories for world model learning. For prompts used to generate these tasks as well as example synthetic tasks generated by GPT-4o, please refer to Section D.3.
We present our results in Figure 4. In Figure 4(b), we find that training with additional world model data steadily improves model’s BoN success rate. This indicates that scaling world model training (with synthetic tasks) enhances the agent’s environment understanding, enabling it to solve novel tasks. However, we did not observe a substantial increase its robustness “robustness” - the agent did not consistently solve these novel tasks across all three trials. We believe this may be due to the stochastic nature of real-world environments, and that increasing world model training data alone cannot ensure the agent to robustly utilize all relevant world knowledge in its policy. We suggest future work should scale both world model and policy training data together, by 1) manually creating a large set of agent tasks with evaluators available for training; and/or 2) developing robust automatic evaluators (Pan et al., 2024) capable of evaluating synthetic tasks generated on-the-fly.
@@ 5.2 Iterating Policy Training in Dyna-Think
We now investigate whether $\pi_{\mathcal{W}}(\theta)$ can be iteratively trained without any supervision from an expert LLM (e.g., GPT-4o was used to critique simulation in Section 3.3). Specifically, we follow STaR (Zelikman et al. (2022), a simple method use for math and reasoning domain) and consider two iterative training loops. Without Evaluation Hint (w/o eval hint), where we 1) first perform rejection sampling using $\pi_{\mathcal{W}}(\theta)$ on the training set; 2) perform policy learning on $\pi_{\mathcal{W}}(\theta)$; and 3) repeat. With Evaluation Hint (w/ eval hint), where for tasks that $\pi_{\mathcal{W}}(\theta)$ fail to solve during step 1, we perform “rationalization” by appending the evaluation configuration to the original instruction (Figure 5(a)), and then perform rejection sampling again on these tasks with evaluation hint. During training and testing, we remove the added hints from the instruction. We report test performance over 5 training iterations.
@@
We present our results in Figure 5(b). We find iterative training without rationalization (w/o eval hint) quickly plateaus; and that training with rationalization (w/ eval hint) outperforms training without rationalization. However, we find no method achieves better than 50% success rate. We believe this is due to the challenging nature of these long-horizon agent tasks. Even when the evaluation configurations are provided as additional hints, we observe that $\pi_{\mathcal{W}}(\theta)$ can only solve a portion of the tasks that it fails to solve otherwise. This indicates that agent tasks, such as computer-use, remains to be a challenging domain for current language models.
@@ 5.3 Quantifying World Model Accuracy
Table 4: Measuring the world model accuracy (Acc) of different policies, and the correlation ($r$) between the average world model accuracy per task and its task success rate.
@@
In this work, we introduced Dyna-Think to integrate world model learning with policy learning in a single $\pi_{\mathcal{W}}(\theta)$. To measure the effectiveness of this world model learning, we now evaluate the World Model Accuracy (Acc.) of different models, and the Pearson Correlation Coefficient ($r$) between the (average) world model accuracy for each task and the task success. To evaluate world model accuracy given an action $a_{i}=\langle\textrm{think}_{i},\textrm{act}_{i}\rangle$ generated by $\pi_{\mathcal{W}}(\theta)$, we 1) prompt GPT-4o to extract the world model simulation text from $\textrm{think}_{i}$ that corresponds to $\textrm{act}_{i}$; and 2) prompt GPT-4o judge whether the extracted simulation is correct given the next state $o_{i+1}$ after executing $a_{i}$. For each model, we calculate this for every turn in each trajectory that terminated within a maximum of 30 steps. Please refer to Section F.1 for prompts used and more details. We present the results in Table 4.
In Table 4, we find that 1) models that achieve a higher success rate also achieve a higher world model accuracy, and that 2) average world model accuracy for each task shows strong correlation with task success, with a minimum correlation of $r=0.32$ across all models. Next, we find that DDT training significantly improved the world model accuracy (16.8% absolute) along with an improved success rate, even though it was trained on the same set of policy data as RFT (see Section 4.3). This shows that combining world-model and policy learning effectively boosts AI agent performance.
@@ 6 Conclusion
We present Dyna-Think, a new thinking framework that synergizes reasoning, acting, and planning by simulating with an internal world model to improve the performance of AI agents. To initialize Dyna-Think, we propose DIT to reconstruct R1’s thinking process and focus on performing compressed world model simulation; and then trains a policy using these reconstructed data. To enhance Dyna-Think, we introduce DDT, a Dyna-style training method to perform both world model training and policy training on a single model. We evaluated our methods on the challenging OSWorld benchmark, and find our models based on Qwen2.5-32B-Instruct reach a similar best-of-n performance compared to R1 (685B), while generating 2x less tokens on average. Our extensive empirical analysis reveals that 1) critique-style world model training is effective for policy improvement; and 2) stronger AI agents show stronger world modeling ability. Our results suggest a promising direction for integrating world model simulation into AI agents to enhance their reasoning, planning, and acting abilities.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have justified claims made in our experiments (Section 4) and discussion sections (Section 5).
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: We provided a limitation section in Appendix A.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: NA.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We provided detailed experimental setups and training details in Section 4.2. We also provided relevant prompts used in our methods in Appendices C, D, E and F
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Yes, we will release our training data and code when accepted.
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We provided detailed training/test dataset statistics in Section 4.1 and optimization details in Section 4.2.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We reported both average performance, best-of-n performance, and standard deviation in our main results in Section 4.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We reported our computation resource in Section 4.2.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We believe our work conforms to NeurIPS code of ethics, and did not cause harms to human subjects or negatively impact the society.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: We have discussed broader impacts of our work in the abstract, conclusion (Section 6), and ethics statement section (Appendix B).
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [Yes]
  44. Justification: Our training data mainly come from OpenAI’s and DeepSeek-R1’s API, which are already moderated for safety guardrails.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We have credited original owners of assets (e.g., data and models).
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We provided detailed training/test dataset statistics in Section 4.1, and information about our trained models in Section 4. We will release these assets when accepted.
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: NA.
  57. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: NA.
  61. 16.
    Declaration of LLM usage
  62. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  63. Answer: [Yes]
  64. Justification: We listed our prompts used in Appendices C, D, E and F.
Appendix A Limitations

### v3

@@ 4.2 Experimental Setup
Since most computer-use benchmarks were not designed for training, the number of tasks available in each domain is often limited. We thus construct a training and test set by 1) manually augmenting existing tasks in each domain to increase its size; 2) construct training/test splits for each domain, and 3) held out two domains (GIMP and Thunderbird) from training to separately measure In-domain (ID) and Out-of-domain (OOD) performance. To augment a task, we follow the principle that 1) action sequence that correctly completes task $A$ does not complete the augmented task $A^{\prime}$; and 2) the augmented task $A^{\prime}$ can still be evaluated using OSWorld’s evaluation scripts after some adjustments. Please refer to Section˜F.2 for more details and example augmented tasks. In Table˜A1 we report the training and testing dataset statistics. For WindowsAgentArena, we directly test on the official tasks as only OSWorld tasks were used in training.
@@
We evaluate all runs using the accessibility-tree mode (i.e., text-only) on both benchmarks, and report task success for ID and OOD tasks. We use the same inference prompts and hyperparameters (e.g., temperature of 1.0) provided by OSWorld, but extend the maximum number of steps per task to 30 in order to measure scaling abilities. To provide a more robust evaluation for these long-horizon tasks, we report the average success rate (Avg) and the Best-of-N success rate (BoN) over 3 runs. Each evaluation is ran with 3 threads, lasting on average 18-24 hours to complete.
@@ 4.3 Main Results
We compare our Dyna-Think framework against 1) training-free methods based on expert LLMs; and 2) related training methods. For training-free methods, we consider prompting LLMs such as o3-mini and DeepSeek-R1 using React. For training-based methods, we consider 1) Reinforcement Finetuning (RFT) which only peforms policy learning, by finetuning on correct rollouts after rejection sampling (Best-of-3); and 2) the vanilla Dyna algorithm (we follow Fang et al. (2025)) which performs world model learning with a separate language model $\mathcal{W}(\mu)$, and then trains the policy using correct rollouts [^1] When rolling out with the learned $\mathcal{W}(\mu)$, correctness of a trajectory is evaluated using an LLM (Fang et al., 2025). We use GPT-4o in this work. collected with both the real $\mathcal{W}$ and the learned $\mathcal{W}(\mu)$. Our DDT method only uses rollouts with $\mathcal{W}$ and performs policy and world model learning on a single model.
For a fair comparison, we use the same set of rollout trajectories (116 in total) to construct world model (116) and policy training data (35 after rejection sampling, a similar data size as Yu et al. (2024b)) for all training methods, whenever possible.
@@
[…]
Table 2: WindowsAgentArena results.
Method Avg Success Rate
[…]
Results
We report our results on OSWorld in Table˜1, and results on WindowsAgentArena on Table˜2. We first compare training-based methods. In Table˜1, we find that DDT training, especially when trained with next-state prediction (DDT($\hat{\mathcal{T}}$)) and critique prediction (DDT($\hat{\mathcal{T}}_{\textrm{critic}}$)) improves upon both RFT and vanilla Dyna in both average success rate and BoN success rate. This indicates that 1) world model training benefits policy training; and 2) directly training $\pi_{\mathcal{W}}(\theta)$ as a world model is more effective than planning with a separately trained $\mathcal{W}(\mu)$. Next, we find that training with critique data showed strong performance compared to training on state-difference (DDT($\hat{\mathcal{T}}_{\Delta}$)) and on next-state prediction (DDT($\hat{\mathcal{T}}$)). We believe this is because these critique data provide a more direct signal for $\pi_{\mathcal{W}}(\theta)$ to improve its world modeling and planning ability during inference (see Section˜5.3 for more quantitative study).
Moreover, we find that training with the next-state prediction objective (DDT($\hat{\mathcal{T}}$)) shows strong performance for OOD tasks, especially on WindowsAgentArena. We believe this is because next-state prediction more effectively enhanced the model’s understanding of the deskptop environments, which is useful for OOD where many states are entirely novel. Finally, compared to best training-free methods such as prompting R1(685B), we find Dyna-Think achieves similar BoN score, while generating 2x less tokens on average and being only 32B in size. These results indicate the many tokens/behaviors during R1-style reasoning may not be necessary, and that focusing on/improving simulation ability is effective at improving agent performance.
@@ Results
[…]
@@ 4.4 Thinking Behavior Analysis
We present the results in Table˜3. First, we find that Direct Distill(R1 no-think) deteriorates significantly compared to Direct Distill(R1), even though they are trained on the same set of trajectories. This indicates that including long-CoT thinking during policy training is benefitial. Next, we find that DIT(R1) achieved similar performance compared to Direct Distill(R1), despite generating 2x less tokens on average. This suggests that the ability to perform reasoning with world-model simulations is the central part in R1-style thinking, underscoring the importance of world modeling in AI agents.
@@ 5.1 Scaling World Model Training in Dyna-Think
Figure 4: Scaling world model learning with synthetic tasks generated by GPT-4o. We use DDT($\hat{\mathcal{T}}_{\textrm{critic}}$) and train from our best model in Table˜3. After world model training, we perform one-round of policy training using the same set of policy learning data.
Since world model learning only requires interaction data (without evaluators for task success) with the real $\mathcal{W}$, we investigate whether scaling world training in Section˜3.3 can further improve policy performance. To test this, we 1) prompt GPT-4o to generate synthetic task instructions for each domain; 2) use $\pi_{\mathcal{W}}(\theta)$ collect rollout trajectories; and 3) construct a world model learning dataset following Section˜3.3. Since no evaluator is required for world modeling data construction, we used all rollout trajectories that terminated within a maximum of 30 steps. This results in a total of 703 additional trajectories for world model learning. For more details please see Section˜E.3.
We present our results in Figure˜4. In Figure˜4(b), we find that training with additional world model data steadily improves model’s BoN success rate. This indicates that scaling world model training (with synthetic tasks) enhances the agent’s environment understanding, enabling it to solve novel tasks. However, we did not observe a substantial increase its “robustness” - the agent did not consistently solve these novel tasks across all three trials. We believe this may be due to the stochastic nature of real-world environments, and that increasing world model training data alone cannot ensure the agent to robustly utilize all relevant world knowledge in its policy. We suggest future work should scale both world model and policy training data together, by 1) manually creating a large set of agent tasks with evaluators available for training; and/or 2) developing robust automatic evaluators (Pan et al., 2024) capable of evaluating synthetic tasks generated on-the-fly.
@@ 5.2 Iterating Policy Training in Dyna-Think
We now investigate whether $\pi_{\mathcal{W}}(\theta)$ can be iteratively trained without any supervision from an expert LLM (e.g., GPT-4o was used to critique simulation in Section˜3.3). Specifically, we follow STaR (Zelikman et al. (2022), a simple method use for math and reasoning domain) and consider two iterative training loops. Without Evaluation Hint (w/o eval hint), where we 1) first perform rejection sampling using $\pi_{\mathcal{W}}(\theta)$ on the training set; 2) perform policy learning on $\pi_{\mathcal{W}}(\theta)$; and 3) repeat. With Evaluation Hint (w/ eval hint), where for tasks that $\pi_{\mathcal{W}}(\theta)$ fail to solve during step 1, we perform “rationalization” by appending the evaluation configuration to the original instruction (Figure˜5(a)), and then perform rejection sampling again on these tasks with evaluation hint. During training and testing, we remove the added hints from the instruction. We report test performance over 5 training iterations.
@@
We present our results in Figure˜5(b). We find iterative training without rationalization (w/o eval hint) quickly plateaus; and that training with rationalization (w/ eval hint) outperforms training without rationalization. However, even when provided with evaluation configurations, we observe that $\pi_{\mathcal{W}}(\theta)$ can only solve a portion of the tasks that it fails to solve otherwise. This indicates that agent tasks, such as computer-use, remains to be a challenging domain for current language models.
@@ 5.3 Quantifying World Model Accuracy
Table 4: Measuring the world model accuracy (Acc) and its correlation ($r$) with task success rate.
@@
In this work, we introduced Dyna-Think to integrate world model learning with policy learning in a single $\pi_{\mathcal{W}}(\theta)$. To measure the effectiveness of this world model learning, we now evaluate the World Model Accuracy (Acc.) of different models, and the Pearson Correlation Coefficient ($r$) between the (average) world model accuracy for each task and the task success. To evaluate world model accuracy given an action $a_{i}=\langle\textrm{think}_{i},\textrm{act}_{i}\rangle$ generated by $\pi_{\mathcal{W}}(\theta)$, we 1) prompt GPT-4o to extract the world model simulation text from $\textrm{think}_{i}$ that corresponds to $\textrm{act}_{i}$; and 2) prompt GPT-4o judge whether the extracted simulation is correct given the next state $o_{i+1}$ after executing $a_{i}$. For each model, we calculate this for every turn in each trajectory that terminated within a maximum of 30 steps. Please refer to Section˜G.1 for prompts used and more details.
In Table˜4, we find that 1) models that achieve a higher success rate also achieve a higher world model accuracy, and that 2) average world model accuracy for each task shows strong correlation with task success, with a minimum correlation of $r=0.32$ across all models. Next, we find that DDT training significantly improved the world model accuracy (16.8% absolute) along with an improved success rate, even though it was trained on the same set of policy data as RFT (see Section˜4.3). This shows that combining world-model and policy learning effectively boosts AI agent performance.
@@ 6 Conclusion
We present Dyna-Think, a new thinking framework that synergizes reasoning, acting, and planning by simulating with an internal world model to improve the performance of AI agents. Dyna-Think consists of two training stages: DIT to initialize a model with simulation ability during reasoning, and DDT for further improvement. We evaluated our methods on OSWorld and WindowsAgentArena, and find our models based on Qwen2.5-32B-Instruct reach a similar best-of-n performance compared to R1(685B), while generating 2x less tokens on average. Our empirical analysis reveals that 1) critique-style world model training is effective for policy improvement; and 2) stronger AI agents show stronger world modeling ability. Our results suggest a promising direction for integrating world model simulation into AI agents to enhance their reasoning, planning, and acting abilities.
Appendix A LLM Usage
This work used LLMs as general-purpose writing assistants to improve the grammar and clarity of the paper. We DO NOT use LLMs to generate research ideas, automate experiments, or analyze results.
Appendix B Limitations